A 3.00-L gas sample at 100.°C and 620. torr contains 93.2% xenon and 6.8% helium by mass. What are the partial pressures of the individual gases?
Answer #1

Let total mass of substance be 100 g

Then mass of Xe = 93.2 g

Then mass of He = 6.8 g

Molar mass of Xe = 131.3 g/mol

Molar mass of He = 4.003 g/mol

n(Xe) = mass of Xe/molar mass of Xe

= 93.2/131.3

= 0.7098

n(He) = mass of He/molar mass of He

= 6.8/4.003

= 1.6987

n(Xe),n1 = 0.7098 mol

n(He),n2 = 1.6987 mol

Total number of mol = n1+n2

= 0.7098 + 1.6987

= 2.4086 mol

Partial pressure of each components are

p(Xe),p1 = (n1*Ptotal)/total mol

= (0.7098 * 620)/2.4086

= 183 torr

p(He),p2 = (n2*Ptotal)/total mol

= (1.6987 * 620)/2.4086

= 437 torr

Answer:

p(Xe) = 183 torr

p(He) = 437 torr